# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 1. $R(x_{k},t_{k})\geq Q_{k}$;
Proof of Claim. Otherwise there is some component $W$ of $\mathcal{O}^{\prime}\setminus\cup_{i}N_{i}$ containing a point $x\in X$ and a point $z\in Z$. We can choose a minimizing geodesic path $\gamma$ in $W$ connecting $x$ and $z$. It is easy to see that the intersection of $\gamma$ with $\partial W$ is empty. For example, one can argue as the following: Suppose the intersection of $\gamma$ with $\partial W$ is not empty. Note that any point $y\in\gamma\cap\partial W$ has a canonical neighborhood $U$ in $\mathcal{O}^{\prime}$. If $U$ is a $4\varepsilon$-cap, since $\delta\ll\varepsilon$ and $y\in\partial W$, the central cross section $S$ of the $4\varepsilon$-neck in the end of the cap, which separates $\mathcal{O}^{\prime}$ (i.e. $\mathcal{O}^{\prime}\setminus S$ is not connected), would be contained in one of the necks $N_{i}$. By the curvature assumption on $x$ and $z$ and the scalar curvature bounds in the definition of $(4\varepsilon,4C)$-canonical neighborhood of cap type, either $x\notin U$ or $z\notin U$, (in fact, since $y\in\partial W$, neither $x$ nor $z$ is contained in $U$; one can also use this fact and the minimizing property of $\gamma$ in $W$ to derive a contradiction,) on the other hand $y$ is in the core of the cap $U$, so $\gamma$ must intersect $S$, hence $\gamma$ must intersect one of the necks $N_{i}$, which is impossible. If $U$ is a $4\varepsilon$-neck, denote by $A$ the boundary component of $W$ which contains $y$, then $\text{diam}~{}A\ll\text{diam}~{}U$. Note that $\gamma$ would enter and exit $U\cap W$ from the same end of $U$. This contradicts the minimizing property of $\gamma$ in $W$. Compare the argument for the first claim in the proof of Lemma 7.7 in [7].
@@
By the definition of $(r,\delta)$-surgical solution, at time $b$ any point $y$ in $\gamma$ of scalar curvature in $[8Cr^{-2},(4C)^{-1}Dh^{-2}]$ has a $(4\varepsilon,4C)$-canonical neighborhood $U$ (in $\mathcal{O}^{\prime}$). $U$ cannot be a component of strictly PIC2, because we have thrown away all such components in $\mathcal{O}(b)$. We will show that $U$ cannot be a $4\varepsilon$-cap either. Otherwise $U=N\cup C$, where $N$ is a $4\varepsilon$-neck, $N\cap C=\emptyset$, $\overline{N}\cap C=\partial C$ and $y\in$ Int $C$. Note that neither $x$ nor $z$ is contained in $U$. Let $\psi:\mathbb{S}^{n-1}/\Gamma\times(-(4\varepsilon)^{-1},(4\varepsilon)^{-1})\rightarrow N$ be the diffeomorphism associated to the neck $N$, and $S=\psi(\mathbb{S}^{n-1}/\Gamma\times\{0\})$. Assume that as the “height” $s\rightarrow(4\varepsilon)^{-1}$ the point in $N$ approaches the end of $U$. Clearly $\gamma$ is not minimizing in $U$, since if $x^{\prime}$ (resp. $z^{\prime}$) is an intersection of $\gamma$ with $S$ between $x$ and $y$ (resp. $y$ and $z$), then $d(x^{\prime},z^{\prime})\ll d(x^{\prime},y)+d(y,z^{\prime})$. The geodesic segment (in $U$) $[x^{\prime}z^{\prime}]$ is not contained in $W$ by the minimality of $\gamma$ in $W$. So $[x^{\prime}z^{\prime}]\cap\partial W\neq\emptyset$. By definition of $W$, the corresponding component of $\partial W$ is a boundary component, denoted by $S_{i}^{+}$, of the closure of some cutoff neck $N_{i}$. Then $d(S_{i}^{+},S)<$diam$(S)$ since $[x^{\prime}z^{\prime}]\cap S_{i}^{+}\neq\emptyset$. It follows that the ratio of the scalar curvature on the neck $N_{i}$ near $S_{i}^{+}$ to that on the neck $N$ near $S$ is very close to 1 (cf. p. 32 in [6] and Proposition A.11 in [67]), and $\text{diam}(S_{i}^{+})\ll\text{diam}(N)$. Then we see that $\psi(\mathbb{S}^{n-1}/\Gamma\times(-(4\varepsilon)^{-1},-(8\varepsilon)^{-1}))\subset W$. This contradicts the minimality of $\gamma$ in $W$. Compare the argument for the second claim in the proof of Lemma 7.7 in [7] and Claim 2 in the proof of Proposition 2.7 in [50]. (By the way, the inequality in lines 7-8 on p. 1225 in [50] should read
@@
Proof.  See the proof of Theorem 8.1.3 in [6] and Proposition 5.2 in [51]. We argue by contradiction. Otherwise there exist $n\geq 3$, $A>0$, $\Lambda>0$, $0<T_{0}<\frac{n-1}{2}$, a sequence $\rho_{k}\rightarrow+\infty$ as $k\rightarrow\infty$, a sequence of open subsets $U_{k}\subset\mathbb{R}^{n}//\Gamma_{k}$, where $\Gamma_{k}$ are finite subgroups of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, such that $B(p_{k},0,\rho_{k}):=B_{\hat{g}_{k}(0)}(p_{k},\rho_{k})\subset U_{k}$ is relatively compact, where $p_{k}:=p_{\Gamma_{k}}$, and $\hat{g}_{k}:=\hat{g}_{\Gamma_{k}}$, and a sequence of Ricci flows $g_{k}(\cdot)$ defined on $U_{k}\times[0,T_{k}]$ with $0<T_{k}\leq T_{0}$, such that
@@
(cf. Lemma 4.5 in [72], Lemma 7.4.1 in [2], Theorem 8.1.2 in [6], Theorem 8.1 in [7], Lemma 7.29 in [56], Proposition 10.5 in [20], Lemma 9.1.1 in [83], and Proposition 5.1 in [51])   Let $\varepsilon$ and $C$ be chosen as before. For any $A>0$ and $\alpha\in(0,\frac{n-1}{2})$ ($n\geq 5$), there exists $\bar{\delta}=\bar{\delta}(A,\alpha)>0$ with the following property. Let $r$ and $\delta$ be two positive numbers with $\delta\leq\bar{\delta}$, and $(\mathcal{O}(\cdot),g(\cdot))$ be a surgical solution of dimension $n$ satisfying the pinching assumption defined on some time interval $[0,b]$, which is an $(r,\delta)$-surgical solution (and in particular, satisfies $(CN)_{r}$ with $(4\varepsilon,4C)$-control) on $[0,b)$. Let $t_{0}\in(0,b)$ be a singular time. Consider the restriction of $(\mathcal{O}(\cdot),g(\cdot))$ to $[t_{0},b]$. Let $p\in\mathcal{O}_{+}(t_{0})$ be the tip of some surgery cap of scale $h=h(r,\delta)$, and let $t_{1}=\min\{b,t_{0}+\alpha h^{2}\}$. Then either
@@
Proof.  Using Lemma 3.6, the proof is almost the same as that of Theorem 8.1.2 in [6] and Theorem 8.1 in [7]. (In fact, when the tip $p$ of a surgery cap is an isolated orbifold singularity, a claim similar to that on p. 95 of [6] can be shown more easily, as in this case the point $p$ can not be in a $\delta$-neck, which is a smooth manifold.) Compare also the second paragraph on p. 69 of [29]. $\Box$
@@
(cf. Proposition 10.9 in [20], Lemma 4.5 in [29], Lemma 5.4 in [51], Proposition 16.1 in [67], and Lemma 5.2 in [72]) Let $(\mathcal{O},g_{0})$ be a compact Riemannian orbifold of dimension $n\geq 5$ with at most isolated singularities and with positive isotropic curvature. Let $\varepsilon$ and $C$ be chosen as before. Then there exist a positive constant $\kappa$ and a positive function $\tilde{\delta}(\cdot)$ with the following property: If we have an $(r,\delta)$-surgical solution $(\mathcal{O}(t),g(t))$, $t\in[0,T]$, to the Ricci flow with $\delta\leq\tilde{\delta}(r)$ starting with $(\mathcal{O},g_{0})$, then for any space-time point $(x_{0},t_{0})$, either the flow is $\kappa$-noncollapsed at $(x_{0},t_{0})$ on all scales less than $\varepsilon$, or $x_{0}$ is contained in a component of strictly PIC2 at time $t_{0}$.
Proof.  The proof is an adaption of that of Proposition 10.9 in [20]; the necessary modifications are as in the proof of in Lemma 4.5 in [29] and Lemma 5.4 in [51], with the help of Propositions 3.7 and 5.4.
@@
Case 2. $r_{0}<\frac{r}{C(n,\theta,\varepsilon)}$. By an argument using orbifold Bishop-Gromov theorem (see [13] and [62]) we may assume w.l.o.g. that there is some point $(x^{\prime},t^{\prime})\in\overline{P(x_{0},t_{0},r_{0},-r_{0}^{2})}$ such that $|\text{Rm}(x^{\prime},t^{\prime})|=r_{0}^{-2}$. (Compare Lemma 10.1.2 in [6], and p.232-233 in [31].) Note that $|\text{Rm}|\leq C(n)R$ for any Riemannian orbifold of dimension $n\geq 5$ with positive isotropic curvature, where $C(n)$ is a positive constant depending only on $n$. So if $C(n,\theta,\varepsilon)$ is large compared to $C(n)$, $(x^{\prime},t^{\prime})$ admits a canonical neighborhood which is not a component of strictly PIC2 (by using our assumption on $x_{0}$). Then using the derivative estimates in the definition of the canonical neighborhood we can get $R(x_{0},t_{0})>r^{-2}$ by choosing $C(n,\theta,\varepsilon)$ sufficiently large. In particular, $(x_{0},t_{0})$ is a center of a $4\varepsilon$-neck or $4\varepsilon$-cap. We consider two subcases.
Subcase 2.1: there is a point $x$ in the connected component containing $x_{0}$ such that $R(x,t_{0})\leq r^{-2}$. In this subcase one can use Proposition 5.4 to find a tube which is diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$, where $\Gamma$ is a finite subgroup of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, such that one end of the tube is adjacent to the $4\varepsilon$-neck or cap centered at $(x_{0},t_{0})$, and $|\text{Rm}|\leq C(n,\theta,\varepsilon)^{2}r^{-2}$ on the (strong) $4\varepsilon$-neck near the other end. By definition of the canonical neighborhood, to bound $\text{vol}_{t_{0}}(B(x_{0},t_{0},r_{0}))/r_{0}^{n}$ from below we only need to control $|\Gamma|$. Then we can reduce this subcase to Case 1; see Step 2 in the proof of Lemma 4.5 in [29].
@@
Proof.   See the proof of Theorem 10.10 in [20]. Suppose the result is not true. Then there exist a sequence of $(\hat{r}_{j},\hat{\delta}_{j})$-surgical solutions to the Ricci flow, $(\mathcal{O}^{(j)}(t),g^{(j)}(t))$, $t\in[0,T_{j}]$, with $\hat{r}_{j}\leq\frac{1}{j}$ and $\hat{\delta}_{j}\leq\text{min}\{\tilde{\delta}(\hat{r}_{j}),\frac{1}{j}\}$, where $\tilde{\delta}(\cdot)$ is the function in Proposition 3.8, (in particular, the solutions satisfy the $(f,\theta)$-pinching assumption and the canonical neighborhood assumptions $(CN)_{\hat{r}_{j}}$ with $(4\varepsilon,4C)$-control,) starting with $(\mathcal{O},g_{0})$, and a sequence of space-time points $(x_{j},t_{j})$ ($x_{j}\in\mathcal{O}^{(j)}(t_{j})$) with $Q_{j}:=R(x_{j},t_{j})\geq(2\hat{r}_{j})^{-2}$ but having no $(2\varepsilon,2C)$-canonical neighborhoods (in particular, $x_{j}$ is not contained in a component which is strictly PIC2 at time $t_{j}$). By Proposition 3.8, the solution $(\mathcal{O}^{(j)}(t),g^{(j)}(t))$ is $\kappa$-noncollapsed on scales less than $\varepsilon$ for some constant $\kappa>0$ independent of $j$. We rescale the Riemannian orbifolds $(\mathcal{O}^{(j)}(t_{j}),g^{(j)}(t_{j}))$ by the factors $Q_{j}$. By Proposition 3.3 the pointed Riemannian orbifolds $(\mathcal{O}^{(j)}(t_{j}),Q_{j}g^{(j)}(t_{j}),x_{j})$ satisfy the so called bounded curvature at bounded distance property. We also have curvature derivative estimates by (using the argument in the proof of) Proposition 10.4 in [20]. By Proposition 4.1 in [56] we see that the pointed Riemannian orbifolds $(\mathcal{O}^{(j)}(t_{j}),Q_{j}g^{(j)}(t_{j}),x_{j})$ subconverge in the pointed Cheeger-Gromov sense to a pointed complete Riemannian orbifold $(\mathcal{O}^{\infty},g^{\infty},x_{\infty})$. By Proposition 6.1 in [29] (see also Proposition 3.18 in [56]) $(\mathcal{O}^{\infty},g^{\infty})$ has bounded curvature.
@@
Using Propositions 2.6, 2.10, 3.7, 5.4 and Lemmas 5.2, 5.3, 5.8 (consult the proof of Lemma 9.2.1 in [6] for the use of Lemma 5.8), we can extend $(\mathcal{O}^{\infty},g^{\infty})$ backwards in time to an orbifold Ricci flow solution $(\mathcal{O}^{\infty},g^{\infty}(t))$ defined on some time interval $(\tau^{*},0]$ (here $\tau^{*}<0$), which has bounded curvature for each $t\in(\tau^{*},0]$. As in Step 6 in the proof of Theorem 7.2 in [20], by using the Harnack inequality in [16] (which extends the work [45]), Lemma 8.3(b) in [71], the $(f,\theta)$-pinching assumption, the canonical neighborhood assumptions $(CN)_{\hat{r}_{j}}$ with $(4\varepsilon,4C)$-control, (the proof of) Proposition 3.3 above, Proposition 4.1 in [56] and Proposition 6.1 in [29] (see also Proposition 3.18 in [56]) we can show that $\tau^{*}=-\infty$. So the parabolical rescalings of the solutions $(\mathcal{O}^{(j)}(t),g^{(j)}(t))$ around $(x_{j},t_{j})$ with factors $Q_{j}$ subconverge to an orbifold ancient $\kappa$-solution satisfying $\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in\text{PIC}$. By using Propositions 2.5, 2.6, 2.9, 5.4 and Lemmas 5.2, 5.3, we see that $(x_{j},t_{j})$ has a $(2\varepsilon,2C)$-canonical neighborhood for $j$ sufficiently large. A contradiction. $\Box$

### v22

@@ 1. $R(x_{k},t_{k})\geq Q_{k}$;
Proof of Claim. Otherwise there is some component $W$ of $\mathcal{O}^{\prime}\setminus\cup_{i}N_{i}$ containing a point $x\in X$ and a point $z\in Z$. We can choose a shortest path $\gamma$ in $W$ connecting $x$ and $z$ w.r.t. the distance induced by the metric $g(b)$; cf. for example Theorem 2.5.23 in [27]. It is not hard to see that the intersection of $\gamma$ with $\partial W$ is empty. For example, one can argue as the following: Suppose the intersection of $\gamma$ with $\partial W$ is not empty. Note that any point $y\in\gamma\cap\partial W$ has a canonical neighborhood $U$ in $\mathcal{O}^{\prime}$. If $U$ is a $4\varepsilon$-cap, by the curvature assumption on $x$ and $z$, the fact that $\frac{1}{2}h^{-2}\leq R(y,b)\leq\frac{3}{2}h^{-2}$, and the scalar curvature bounds in the definition of $(4\varepsilon,4C)$-canonical neighborhood of the cap type, neither $x$ nor $z$ is contained in $U$. Write $U=N\cup C$, where $N$ is a $4\varepsilon$-neck in the end of the cap $U$, $N\cap C=\emptyset$, $\overline{N}\cap C=\partial C$ and $y\in\text{Int}\hskip 1.42262ptC$. Let $N_{i_{0}}$ be the $\delta$-neck whose closure contains the point $y$. We may assume that the point $y$ is near the negative end of the neck $N_{i_{0}}$ (w.r.t. the height function). First observe that $C\cap N_{i_{0}}\neq\emptyset$. Otherwise any neighborhood of $y$ (in $\mathcal{O}^{\prime}$) must contain a point in $N_{i_{0}}$, which is not in $C$. This is impossible, as $\text{Int}\hskip 1.42262ptC$ is an open subset of $\mathcal{O}^{\prime}$. Then it is easy to see that the supremum of the height function on the neck $N_{i_{0}}$ restricted to $C\cap N_{i_{0}}$ is attainable, and is attained at a point in $\partial C$ instead of $\text{Int}\hskip 1.42262ptC$ by an argument as before using again the fact that $\text{Int}\hskip 1.42262ptC$ is an open subset of $\mathcal{O}^{\prime}$. In particular, $\partial C\cap N_{i_{0}}\neq\emptyset$. Note that $\partial C$ is a central cross-section of a $4\varepsilon$-neck by definition of $4\varepsilon$-cap, and $\delta\ll\varepsilon$. From the proof of Proposition A.4 we see that the central cross-section $S$ of the neck $N$ is contained in $N_{i_{0}}$. But this is impossible, since on the one hand, as $S$ separates $\mathcal{O}^{\prime}$ (i.e. $\mathcal{O}^{\prime}\setminus S$ is not connected), the path $\gamma$ must intersect $S$; on the other hand, $\gamma$ can not intersect $N_{i_{0}}$, in particular, $\gamma$ can not intersect $S$. (Alternatively, one can also argue as follows. Observe that for any $i$, the middle nine-tenths of the neck $N$ can not intersect $N_{i}$, for otherwise suppose that the middle nine-tenths of the neck $N$ intersects some $N_{i}$, then by the proof of Proposition A.4 there would exist a cross-section of $N$, say $S^{\prime}$, which is contained in $N_{i}$. But this is impossible as seen above. So the middle nine-tenths of the neck $N$ is contained in $W$. But this is also impossible, as the path $\gamma$ must intersect the central cross-section of the neck $N$ at least twice, $\gamma$ can not be minimizing in the situation that the middle nine-tenths of the neck $N$ is contained in $W$, which is a contradiction.) If $U$ is a $4\varepsilon$-neck, denote by $A$ the boundary component of $W$ which contains $y$, then $\text{diam}~{}A\ll\text{diam}~{}U$. Note that $\gamma$ would enter and exit $U\cap W$ from the same end of $U$. Again this contradicts the minimizing property of $\gamma$ in $W$. Compare the argument for the first claim in the proof of Lemma 7.7 in [7].
@@
By the definition of $(r,\delta)$-surgical solution, at time $b$ any point $y$ in $\gamma$ of scalar curvature in $[8Cr^{-2},(4C)^{-1}Dh^{-2}]$ has a $(4\varepsilon,4C)$-canonical neighborhood $U$ (in $\mathcal{O}^{\prime}$). $U$ cannot be a component of strictly PIC2, because we have thrown away all such components in $\mathcal{O}(b)$. We will show that $U$ cannot be a $4\varepsilon$-cap either. Otherwise similarly as above write $U=N\cup C$, where $N$ is a $4\varepsilon$-neck, $N\cap C=\emptyset$, $\overline{N}\cap C=\partial C$ and $y\in\text{Int}\hskip 1.42262ptC$. Note again that neither $x$ nor $z$ is contained in $U$. Let $\psi:\mathbb{S}^{n-1}/\Gamma\times(-(4\varepsilon)^{-1},(4\varepsilon)^{-1})\rightarrow N$ be the diffeomorphism associated to the neck $N$, and $S=\psi(\mathbb{S}^{n-1}/\Gamma\times\{0\})$. Assume that as the “height” $s\rightarrow(4\varepsilon)^{-1}$ the point in $N$ approaches the end of $U$. Clearly $\gamma$ is not minimizing in $U$, since if $x^{\prime}$ (resp. $z^{\prime}$) is an intersection of $\gamma$ with $S$ between $x$ and $y$ (resp. $y$ and $z$), then $d(x^{\prime},z^{\prime})\ll d(x^{\prime},y)+d(y,z^{\prime})$. The geodesic segment (in $U$) $[x^{\prime}z^{\prime}]$ is not contained in $W$ by the minimality of $\gamma$ in $W$. So $[x^{\prime}z^{\prime}]\cap\partial W\neq\emptyset$. By definition of $W$, the corresponding component of $\partial W$ is a boundary component, denoted by $S_{i}^{+}$, of the closure of some cutoff neck $N_{i}$. Then $d(S_{i}^{+},S)<\text{diam}(S)$ since $[x^{\prime}z^{\prime}]\cap S_{i}^{+}\neq\emptyset$. It follows that the ratio of the scalar curvature on the neck $N_{i}$ near $S_{i}^{+}$ to that on the neck $N$ near $S$ is very close to 1 (cf. p. 32 in [6] and Proposition A.11 in [70]), and $\text{diam}(S_{i}^{+})\ll\text{diam}(N)$. Then we see that $\psi(\mathbb{S}^{n-1}/\Gamma\times(-(4\varepsilon)^{-1},-(8\varepsilon)^{-1}))\subset W$. This contradicts the minimality of $\gamma$ in $W$. Compare the argument for the second claim in the proof of Lemma 7.7 in [7] and Claim 2 in the proof of Proposition 2.7 in [53]. (By the way, the inequality in lines 7-8 on p. 1225 in [53] should read
@@
Proof.  See the proof of Theorem 8.1.3 in [6] and Proposition 5.2 in [54]. We argue by contradiction. Otherwise there exist numbers $n\geq 3$, $A>0$, $\Lambda>0$, $0<T_{0}<\frac{n-1}{2}$, a sequence $\rho_{k}\rightarrow+\infty$ as $k\rightarrow\infty$, a sequence of open subsets $U_{k}\subset\mathbb{R}^{n}//\Gamma_{k}$, where $\Gamma_{k}$ are finite subgroups of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, such that $B(p_{k},0,\rho_{k}):=B_{\hat{g}_{k}(0)}(p_{k},\rho_{k})\subset U_{k}$ is relatively compact, where $p_{k}:=p_{\Gamma_{k}}$, and $\hat{g}_{k}:=\hat{g}_{\Gamma_{k}}$, and a sequence of Ricci flows $g_{k}(\cdot)$ defined on $U_{k}\times[0,T_{k}]$ with $0<T_{k}\leq T_{0}$, such that
@@
(cf. Lemma 4.5 in [76], Theorem 8.1.2 in [6], Theorem 8.1 in [7], Lemma 7.3.6 in [29], Proposition 10.5 in [21], Lemma 9.1.1 in [90], and Proposition 5.1 in [54])   Fix $n\geq 5$, and let $\varepsilon$ and $C$ be chosen as before. For any $A>0$ and $\alpha\in(0,\frac{n-1}{2})$, there exists $\bar{\delta}=\bar{\delta}(A,\alpha)>0$ with the following property. Let $r$ and $\delta$ be two positive numbers with $\delta\leq\bar{\delta}$, and $(\mathcal{O}(\cdot),g(\cdot))$ be a complete surgical solution of dimension $n$ with positive isotropic curvature defined on some time interval $[0,b]$, which is an $(r,\delta)$-surgical solution (and in particular, satisfies $(CN)_{r}$ with $(4\varepsilon,4C)$-control) on $[0,b)$. Let $t_{0}\in(0,b)$ be a singular time. Consider the restriction of $(\mathcal{O}(\cdot),g(\cdot))$ to $[t_{0},b]$. Let $p\in\mathcal{O}_{+}(t_{0})$ be the tip of some surgery cap of scale $h=h(r,\delta)$, and let $t_{1}=\min\{b,t_{0}+\alpha h^{2}\}$. Then either
@@
Proof.  Using Lemma 3.6 and the fact that $|\text{Rm}|\leq C(n)R$ for any Riemannian orbifold of dimension $n\geq 5$ with positive isotropic curvature, where $C(n)$ is a positive constant depending only on $n$, the proof is almost the same as that of Theorem 8.1.2 in [6] and Theorem 8.1 in [7]. Note that here we do not need to use the pinching assumption in the definition of $(r,\delta)$-surgical solution. $\Box$
@@
(cf. Proposition 10.9 in [21], Lemma 4.5 in [31], Lemma 5.4 in [54], Proposition 16.1 in [70], and Lemma 5.2 in [76]) Let $(\mathcal{O},g_{0})$ be a compact Riemannian orbifold of dimension $n\geq 5$ with at most isolated singularities and with positive isotropic curvature. Let $\varepsilon$ and $C$ be chosen as before. Then there exists a positive constant $\kappa$ and a positive function $\tilde{\delta}(\cdot)$ with the following property: If we have an $(r,\delta)$-surgical solution $(\mathcal{O}(t),g(t))$, $t\in[0,T]$, to the Ricci flow with $\delta\leq\tilde{\delta}(r)$ starting with $(\mathcal{O},g_{0})$, then for any space-time point $(x_{0},t_{0})$, either the flow is $\kappa$-noncollapsed at $(x_{0},t_{0})$ on all scales less than $\varepsilon$, or $x_{0}$ is contained in a component of strictly PIC2 at time $t_{0}$.
Proof.  The proof is an adaption of that of Proposition 10.9 in [21]; the necessary modifications are as in the proof of in Lemma 4.5 in [31] and Lemma 5.4 in [54], with the help of Propositions 3.7 and A.4.
@@
Case 2. $r_{0}<\frac{r}{C(n,\theta,\varepsilon)}$. By an argument using orbifold Bishop-Gromov theorem (see [13] and [65]) we may assume w.l.o.g. that there is some point $(x^{\prime},t^{\prime})\in\overline{P(x_{0},t_{0},r_{0},-r_{0}^{2})}$ such that $|\text{Rm}(x^{\prime},t^{\prime})|=r_{0}^{-2}$. (Compare Lemma 10.1.2 in [6], and p.232-233 in [33].) As before, note that $|\text{Rm}|\leq C(n)R$ for any Riemannian orbifold of dimension $n\geq 5$ with positive isotropic curvature, where $C(n)$ is a positive constant depending only on $n$. So if $C(n,\theta,\varepsilon)$ is large compared to $C(n)$, $(x^{\prime},t^{\prime})$ admits a canonical neighborhood which is not a component of strictly PIC2 (by using our assumption on $x_{0}$). Then using the derivative estimates in the definition of the canonical neighborhood we can get $R(x_{0},t_{0})>r^{-2}$ by choosing $C(n,\theta,\varepsilon)$ sufficiently large. In particular, $(x_{0},t_{0})$ is a center of a $4\varepsilon$-neck or $4\varepsilon$-cap. We consider two subcases.
Subcase 2.1: there is a point $x$ in the connected component containing $x_{0}$ such that $R(x,t_{0})\leq r^{-2}$. In this subcase one can use Proposition A.4 and Lemma 3.1 to find a tube which is diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$, where $\Gamma$ is a finite subgroup of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, such that one end of the tube is adjacent to the $4\varepsilon$-neck or cap centered at $(x_{0},t_{0})$, and $|\text{Rm}|\leq C(n,\theta,\varepsilon)^{2}r^{-2}$ on the (strong) $4\varepsilon$-neck near the other end. By definition of the canonical neighborhood, to bound $\text{vol}_{t_{0}}(B(x_{0},t_{0},r_{0}))/r_{0}^{n}$ from below we only need to control $|\Gamma|$. Then we can reduce this subcase to Case 1; see Step 2 in the proof of Lemma 4.5 in [31].
@@
Proof.   See the proof of Theorem 10.10 in [21]. Suppose the result is not true. Then there exists a sequence of $(\hat{r}_{j},\hat{\delta}_{j})$-surgical solutions to the Ricci flow, $(\mathcal{O}^{(j)}(t),g^{(j)}(t))$, $t\in[0,T_{j}]$, with $\hat{r}_{j}\leq\frac{1}{j}$ and $\hat{\delta}_{j}\leq\text{min}\{\tilde{\delta}(\hat{r}_{j}),\frac{1}{j}\}$, where $\tilde{\delta}(\cdot)$ is the function in Proposition 3.8, (in particular, the solutions satisfy the $(f,\theta)$-pinching assumption and the canonical neighborhood assumptions $(CN)_{\hat{r}_{j}}$ with $(4\varepsilon,4C)$-control,) starting with $(\mathcal{O},g_{0})$, and a sequence of space-time points $(x_{j},t_{j})$ ($x_{j}\in\mathcal{O}^{(j)}(t_{j})$) with $Q_{j}:=R(x_{j},t_{j})\geq(2\hat{r}_{j})^{-2}$ but having no $(2\varepsilon,2C)$-canonical neighborhoods (in particular, $x_{j}$ is not contained in a component which is strictly PIC2 at time $t_{j}$). By Proposition 3.8, the solution $(\mathcal{O}^{(j)}(t),g^{(j)}(t))$ is $\kappa$-noncollapsed on scales less than $\varepsilon$ for some constant $\kappa>0$ independent of $j$. We rescale the Riemannian orbifolds $(\mathcal{O}^{(j)}(t_{j}),g^{(j)}(t_{j}))$ by the factors $Q_{j}$. By Proposition 3.3 the pointed Riemannian orbifolds $(\mathcal{O}^{(j)}(t_{j}),Q_{j}g^{(j)}(t_{j}),x_{j})$ satisfy the so called bounded curvature at bounded distance property. We also have curvature derivative estimates by (using the argument in the proof of) Proposition 10.4 in [21]. By Proposition 4.1 in [59] (see also Theorem 3.5 in [65]) we see that the pointed Riemannian orbifolds $(\mathcal{O}^{(j)}(t_{j}),Q_{j}g^{(j)}(t_{j}),x_{j})$ subconverge in the pointed Cheeger-Gromov sense to a pointed complete Riemannian orbifold $(\mathcal{O}^{\infty},g^{\infty},x_{\infty})$. By Proposition 6.1 in [31] (see also Proposition 3.18 in [59]) $(\mathcal{O}^{\infty},g^{\infty})$ has bounded curvature.
@@
Using Propositions 2.6, 2.10, 3.7, A.4 and Lemmas A.2, A.3, A.8 (consult the proof of Lemma 9.2.1 in [6] for the use of Lemma A.8), we can extend $(\mathcal{O}^{\infty},g^{\infty})$ backwards in time to an orbifold Ricci flow solution $(\mathcal{O}^{\infty},g^{\infty}(t))$ defined on some time interval $(\tau^{*},0]$ (here $\tau^{*}<0$), which has bounded curvature for each $t\in(\tau^{*},0]$. As in Step 6 in the proof of Theorem 7.2 in [21], by using the Harnack inequality in [17] (which extends the work [48]), Lemma 8.3(b) in [75], the $(f,\theta)$-pinching assumption, the canonical neighborhood assumptions $(CN)_{\hat{r}_{j}}$ with $(4\varepsilon,4C)$-control, (the proof of) Proposition 3.3 above, Proposition 4.1 in [59] and Proposition 6.1 in [31] (see also Proposition 3.18 in [59]) we can show that $\tau^{*}=-\infty$. So the parabolical rescalings of the solutions $(\mathcal{O}^{(j)}(t),g^{(j)}(t))$ around $(x_{j},t_{j})$ with factors $Q_{j}$ subconverge to an orbifold ancient $\kappa$-solution satisfying $\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in\text{PIC}$. By using Propositions 2.5, 2.6, 2.9, A.4 and Lemmas A.2, A.3, we see that $(x_{j},t_{j})$ has a $(2\varepsilon,2C)$-canonical neighborhood for $j$ sufficiently large. A contradiction. $\Box$